OpenStudy (anonymous):

2squareroot3+5squareroot2 over 3squareroot3 -2squareroot2

OpenStudy (anonymous):

\[2\sqrt{3}+5\sqrt{2} \over 3\sqrt{3}-2\sqrt{2}\]

OpenStudy (primeralph):

Rationalize it.

OpenStudy (anonymous):

Idk how to do that lol

OpenStudy (primeralph):

@robz Please help here. I got to answer something else.

OpenStudy (primeralph):